OpenStudy (anonymous):

Let f(x) =4x-1 and g(x) = 2x^2 + 3 (fg)(x) , f/g(x), g/f(x)

Nnesha (nnesha):

(fg)(x) multiply both function

Nnesha (nnesha):

\[\huge\rm (fg)(x) = (4x-1)(2x^2+3)\] apply foil method :-)

OpenStudy (anonymous):

so 8x2 + 12 - 2x2 - 3? @Nnesha

Nnesha (nnesha):

when you multiply variables you should add their exponent so remember x <----there is invisible exponent which is ?

Nnesha (nnesha):

\[\huge\rm x \times x^2 = x^?\]

OpenStudy (anonymous):

1... so it should be 8x3?

Nnesha (nnesha):

yep!

OpenStudy (anonymous):

now do you just simplify and thats your answer?

Nnesha (nnesha):

look over ur answer one more time 4x times 3 = ?? :-)

OpenStudy (anonymous):

12x whoops

OpenStudy (anonymous):

so 8x3 + 12x - 2x2 - 3

Nnesha (nnesha):

yep :-) so \[\huge\rm 8x^3 +12x -2x^2 -3\] :-)

OpenStudy (anonymous):

so we combine and thats the final answer ??

Nnesha (nnesha):

can you combine anything ?? are there any like terms to combine ?? :-)

OpenStudy (anonymous):

well not all the x's are the same.. so maybe not

Nnesha (nnesha):

yes so that's it :-)

OpenStudy (anonymous):

hm. seems simple enough

Nnesha (nnesha):

yeah :-)

OpenStudy (anonymous):

can you help with just one more that is a bit harder

Nnesha (nnesha):

sure ! what about f/g and g/f can you divide that ?? or do you need help ?

OpenStudy (anonymous):

oh i completely forgot about those! okay well i know you start off with 4x - 1/2x2 + 3 and 2x2+3/4x-1 but I'm not sure where to go from there

Nnesha (nnesha):

okay let's do first one \[\huge\rm \frac{ 4x-1 }{ 2x^2 +3 }\] can you factor top or bottom part ?

OpenStudy (anonymous):

i don't think so... @Nnesha

Nnesha (nnesha):

yes there isn't any common factor so done!

Nnesha (nnesha):

\[\huge\rm \frac{ 2x^2+3 }{ 4x -1 }\] hmm for this one u can use long division ?

Nnesha (nnesha):

have u learned "long division " topic

Nnesha (nnesha):

??*